OpenStudy (anonymous):

Use = 3.14 to estimate the circumference of the circle to the nearest hundredth. (1 point) A)34.54 cm B)17.27 cm C)5.5 cm D)33 cm

OpenStudy (anonymous):



OpenStudy (anonymous):

Circumference of a circle: \[\pi \times 2 \times radius\]or \[\pi \times diameter\]

OpenStudy (anonymous):

\[ 3.14 \times 11 = ?\]

OpenStudy (anonymous):

Use a calculator. What is your answer?

OpenStudy (anonymous):

34.54

OpenStudy (anonymous):

Perfect. (:

OpenStudy (anonymous):

thank you very much you should be a teacher

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

No problem. After you practice using the formulas they'll eventually stick, trust me. (;

OpenStudy (anonymous):

thx